# Unifews: Unified Entry-Wise Sparsification for Efficient Graph Neural Network

### v1

@@ 3.1. Graph Neural Network
3.2. Graph Smoothing
A broad scope of canonical GNN learning processes including both iterative and decoupled architectures can be characterized by a graph smoothing process with a unified optimization objective (Ma et al., 2021; Zhu et al., 2021). We employ this framework as an approach to study GNN by examining the overall learning objective:
Definition 3.1 (Graph Laplacian Smoothing (Ma et al., 2021)).
(4) $\bm{p}^{\ast}=\operatorname*{arg\,min}_{\bm{p}}\mathcal{L},\quad\mathcal{L}=\|\bm{p}-\bm{x}\|^{2}+c\cdot\bm{p}^{\top}\bm{L}\bm{p},\vspace{-4pt}$
[…]
In Eq. 4, $\bm{L}$ is the general Laplacian matrix, as normalization only causes a difference in coefficient. The first term of objective $\mathcal{L}$ reflects the closeness to the input signal, which is generally node attributes representing their identity. The second term is associated with the graph structure, acting as a regularization that constrains the representation values of connected node pairs to be similar. In other words, the implication of Eq. 4 optimization process is to strike a balance between the attribute and structural information of the graph as the produced embedding. The regularization coefficient $c$ is employed to control the relative strength between the two terms. The closed-form solution $\bm{p}^{\ast}$ of Eq. 4 can be inferred when the derivative of optimization goal equals 00:
(5) $\frac{\partial\mathcal{L}}{\partial\bm{p}}=2(\bm{p}-\bm{x})+2c\bm{L}\bm{p}=0\;\Rightarrow\;\bm{p}^{\ast}=(\bm{I}+c\bm{L})^{-1}\bm{x}.$
More generally, when there are $f$ input vectors, i.e., the input matrix is $\bm{X}\in\mathbb{R}^{n\times f}$, then the matrix form of graph Laplacian smoothing corresponding to Eq. 4 is:
(6) $\bm{P}^{\ast}=\operatorname*{arg\,min}_{\bm{P}}\|\bm{P}-\bm{X}\|_{F}^{2}+c\cdot\operatorname{tr}(\bm{P}^{\top}\bm{L}\bm{P}),\vspace{-4pt}$
where $\|\cdot\|_{F}$ is the matrix Frobenius norm and $\bm{P}^{\ast}=(\bm{I}+c\bm{L})^{-1}\bm{X}$.
Although Eq. 5 derives the converged solution of the graph smoothing problem, it is prohibitive to directly acquire it due to the inverse calculation on the large graph-scale matrix. Hence, GNN models employ an iterative approach to learn the representation by stacking multiple layers as described in Section 3.1. Their architectural designs differentiate from each other regarding the exact smoothing settings. For example, GCN convolution corresponds to $c=1$, while GAT embraces a variable coefficient $c[u,v]$ related with the attention of each node pair $(u,v)$ (Ma et al., 2021).
(a) Propagation Personalization
(b) Architecture Compression
(c) Unifews on Iterative GNN
Figure 2. Comparison of GNN learning pipelines between conventional simplification techniques and our Unifews framework. (a) The approach of personalized propagation iteratively simplifies the node-dependent graph diffusion but is only applicable to decoupled GNNs. (b) Joint model compression can sparsify both graph and weight, whereas the same diffusion is uniformly utilized across all layers. (c) In contrast, our Unifews framework implements on-the-fly unified sparsification of both graph and weights for each layer. The adaptive entry-wise scheme enjoys fine-grained control and improved sparsity.
4. Unifews Sparsification
This section presents the Unifews framework by respectively developing its application to the two categories of GNN architectures. Initially, we investigate the decoupled design in Section 4.1, where the graph propagation stage is isolated, offering an ideal context featuring the graph sparsification approach in Unifews. Weight sparsification is then incorporated into the iterative model transformation, as described in Section 4.2.
4.1. Applying to Decoupled GNN
[…]
Key Intuition: Entries in Graph Computation. Conventional graph sparsification methods for GNNs only provide fixed and uniform graph-level adjustments through the entire learning process. This lack of flexibility hinders their performance when employing to the recurrent update design with multiple GNN layers. Contrarily, node-wise models aim to personalize the graph-based propagation, but come with additional explicit calculations and impaired runtime efficiency. We are hence motivated to design a sparsification approach that (1) enables modifications in a fine-grained manner to further simplify the computation; and (2) can be seamlessly integrated into the operation with negligible overhead.
The core concept of our Unifews sparsification is to evaluate and prune elements at the entry level. As illustrated in Figure 0(a), in a single hop of GNN diffusion, each graph connection carries a propagation message from the source node to target. Presenting the procedure by the diffusion matrix $\bm{T}$, each entry $\bm{T}[u,v]$, i.e., the element in the row $u$ and column $v$ of the matrix $\bm{T}$, corresponds to a graph edge, and its value represents the weight of edge propagation. Without loss of generality, we express this entry-wise graph propagation by rewriting Eq. 2 for the embedding on node $u$ as:
(7) $\bm{p}_{(l+1)}[u]=\sum_{v\in\mathcal{N}(u)}\tau_{(l)}[u,v]=\sum_{v\in\mathcal{N}(u)}\bm{T}_{(l)}[u,v]\cdot\bm{p}_{(l)}[v],$
where $\bm{p}[v]$ is the individual embedding value of node $v$, and the product $\tau_{(l)}[u,v]=\bm{T}_{(l)}[u,v]\bm{p}_{(l)}[v]$, i.e., the propagation message from node $v$ to $u$, is regarded as an entry. Eq. 7 presents an entry-centric view of graph propagation, where messages are first calculated for each edge, and subsequently aggregated to the center node. Similar to propagation personalization approaches, we also denote the graph diffusion $\bm{T}_{(l)}$ in a layer-specific fashion.
[…]
Importance-based Pruning. The example in Figure 0(a) suggests that, edges carrying propagation messages exert varying impacts on neighboring nodes, influenced by the graph topology. Messages with minor significance are usually considered redundant and can be eliminated to facilitate sparsity. From the perspective of matrix computation, this is achieved by omitting the particular message $\tau[u,v]$ in Eq. 7 aggregation for the current layer based on an importance indicator such as its magnitude.
However, the message entry $\tau[u,v]$ is dependent on the variable embedding $\bm{p}[v]$. To rigidly record the pruning information for future use, we opt to utilize the diffusion matrix. This approach equates the removal of a message to zeroing out the corresponding entry $\bm{T}[u,v]$. We hence propose the entry-wise graph sparsification scheme that alters the graph diffusion entry as:
(8) $\hat{\bm{T}}[u,v]=\operatorname{thr}_{\delta_{a}}\big{(}\tau[u,v]\big{)}\cdot\bm{T}[u,v],$
where $\delta_{a}$ denotes the sparsification threshold parameter, and the pruning function with an arbitrary threshold $\delta$ is:
$\operatorname{thr}_{\delta}(x)=\left\{\begin{array}[]{lc}1,&|x|>\delta\\ 0,&|x|\leq\delta\end{array}\right..$
In certain practical scenarios, dropping entries during propagation calculation by Eq. 8 may be difficult due to the fixed implementation of matrix multiplication operators. However, as the values of $\bm{T}$ and $\bm{p}$ are known in prior, the edges for diffusion can be equivalently processed beforehand with a node-wise threshold:
(9) $\hat{\bm{T}}[u,v]=\operatorname{thr}_{\delta^{\prime}_{a}}\big{(}\bm{T}[u,v]\big{)}\cdot\bm{T}[u,v],\quad\delta^{\prime}_{a}=\delta_{a}/|\bm{p}[v]|.$
Sparsification by Eq. 8 and Eq. 9 can be applied interchangeably depending on the specific implementation. It has two concurrent effects: for graph computation, messages with small magnitudes are prevented from propagating to neighboring nodes and composing the output embedding; for graph topology, corresponding edges are removed from the graph diffusion process. An example on the real-word graph cora is shown in Figure 2(a), which demonstrates that the distribution of message values is correlated with the diffusion entries. Hence, message magnitude is effective in representing edge importance and determining entry removal.
Algorithm 1 Unifews on Decoupled Propagation
Graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$, diffusion $\bm{T}_{(l)}$, attribute $\bm{X}$, propagation hop $L$, graph sparsification threshold $\delta_{a}$pproximate embedding $\hat{\bm{P}}_{(L)}$
[…]
Accumulation across Layers. Thanks to the adaptive property of the entry-wise pruning scheme, we are able to perform fine-grained and gradual sparsification across propagation layers. Intuitively, a message deemed of minor significance in the current propagation is unlikely to propagate further and influence more distant nodes in subsequent layers. Therefore, if an edge is pruned in layer $l_{0}$, it can be excluded from all following layers $l\geq l_{0}$. By this means, Unifews ensures progressively increasing sparsity throughout the iterative updates of GNN layers.
As depicted in Figure 1(c), Unifews sparsification iteratively applies Eq. 8 to each GNN layer, and the pruned diffusion matrix $\hat{\bm{T}}_{(l)}$ is inherited to the next layer for further edge removal. Denote the edge set corresponding to $\hat{\bm{T}}_{(l)}$ as ^$\hat{\mathcal{E}}_{(l)}$, there is ^$\hat{\mathcal{E}}_{(l)}\subseteq\hat{\mathcal{E}}_{(l-1)}\subseteq\cdots\subseteq\hat{\mathcal{E}}_{(0)}\subseteq\mathcal{E}$, which indicates a diminishing number of propagation operations for deeper layers in the network. Consequently, Eq. 7 is modified as the following to represent entry-level diffusion on the sparsified graph:
(10) $\hat{\bm{p}}_{(l+1)}[u]=\sum_{v\in\mathcal{N}(u)}\hat{\bm{T}}_{(l)}[u,v]\bm{p}_{(l)}[v]=\sum_{v\in\hat{\mathcal{N}}_{(l)}(u)}\tau_{(l)}[u,v],\vspace{-2pt}$
where the neighborhood composed by remaining connections is $\hat{\mathcal{N}}_{(l)}(u)=\{v|(u,v)\in\hat{\mathcal{E}}_{(l)}\}$. For each layer, denote the number of removed edges as $q_{a}=m-|\hat{\mathcal{E}}_{(l)}|$, the graph sparsity is $\eta_{a}=q_{a}/m$. Thence the computation complexity for layer propagation is reduced from $O(m)$ to $O\big{(}(1-\eta_{a})m\big{)}$.
We summarize the Unifews graph sparsification process applied to the propagation process in decoupled GNNs in Algorithm 1. For the matrix with feature dimension $f$, the threshold function Eq. 9 is performed by replacing the absolute value $|\bm{p}[v]|$ with the norm of corresponding node representation vector $\|\bm{P}[v]\|$ across all feature dimensions. Note that we adopt skip connections to each layer in the algorithm by initializing $\bm{P}_{(l+1)}$ explicitly with the value in the previous hop, in order to preserve at least one meaningful value for each node embedding in case all connected edges are pruned.
[…]
@@ 4.1. Applying to Decoupled GNN
Comparative Discussion. We highlight three advantages of the entry-wise approach in the context of GNN learning. Firstly, realistic implementations for graph propagation and feature transformation are also conducted in an entry-centric fashion, by first computing the values of all entries and then performing aggregation for each node to acquire the feature vector. Therefore, entry-level operations can be naturally inserted into the computation on the fly without additional overhead. Secondly, different from previous sparsification strategies introduced in Section 3 such as node-dependent schemes on the product vector $\bm{p}$ or edge pruning only considering $\bm{T}$, our entry-wise sparsification operates on the exact messages that are being passed during propagation, which inherently carries the information that can be utilized for pruning. It hence offers more adaptive and fine-grained control on graph connections for better efficiency-efficacy trade-off. Lastly, the incremental pruning across layers is capable of promoting greater graph sparsity and filtering only important messages during learning, thereby alleviating intrinsic drawbacks of the GNN architecture including over-smoothing and neighborhood explosion as analyzed in Section 5.
Moreover, the feature transformation stage Eq. 3 in decoupled GNNs is feasible to the weight pruning described in the following section, which is similar to classical irregular pruning with magnitude-based thresholds for MLP networks (Han et al., 2015; Srinivas and Babu, 2015).
4.2. Applying to Iterative GNN
Compared to decoupled designs, graph propagation and feature transformation in iterative GNN models are tightly integrated in each layer. Traditionally, this poses a challenge to GNN sparsification methods depicted in Figure 1(b), as specialized schemes are necessary for simultaneously modifying the two components without impairing the performance. On the contrary, our Unifews approach takes advantage of the opportunity for jointly sparsifying the model towards a win-win situation in graph learning. Its framework within and between iterative GNN layers is shown in Figures 0(b) and 1(c), respectively.
Algorithm 2 Unifews on Iterative GNN
Graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$, diffusion $\bm{T}_{(l)}$, attribute $\bm{X}$, network layer $L$, graph and weight sparsification thresholds $\delta_{a},\delta_{w}$pproximate representation $\hat{\bm{H}}_{(L)}$
[…]
The sparsification of iterative Unifews for the entire message-passing scheme is presented in Algorithm 2. Its graph propagation stage is identical to Algorithm 1 for each hop, except that the embedding is initialized by previous representation $\hat{\bm{H}}$ and the corresponding entry is $\bm{\tau}[u,v]=\bm{T}[u,v]\bm{H}[v]$.
After obtaining embedding $\hat{\bm{P}}_{(l)}$ for the current layer, feature transformation is performed by multiplying the weight matrix. We similarly rewrite Eq. 1 with respect to each column $i$, i.e., the $i$-th feature dimension of the representation matrix:
(11) $\small\bm{H}_{(l+1)}[:,i]=\sigma\Big{(}\sum_{j=1}^{f}\bm{\omega}_{(l)}[j,i]\Big{)}=\sigma\Big{(}\sum_{j=1}^{f}\bm{W}_{(l)}[j,i]\cdot\bm{P}_{(l)}[:,j]\Big{)},\vspace{-3pt}$
(12) $\hat{\bm{W}}[j,i]=\operatorname{thr}_{\delta^{\prime}_{w}}\big{(}\bm{W}[j,i]\big{)}\cdot\bm{W}[j,i],\quad\delta^{\prime}_{w}=\delta_{w}/\|\bm{P}[:,j]\|.$
Given the shapes of matrices $\bm{P},\bm{H}\in\mathbb{R}^{n\times f}$ and $\bm{W}\in\mathbb{R}^{f\times f}$, let the pruning ratio of weight matrix be $\eta_{w}$. The sparsification scheme at least reduces complexity to $O\big{(}(1-\eta_{w})nf^{2}\big{)}$. The favorable merit of joint graph and weight sparsification by Unifews is that, both the propagation result $\bm{P}$ and the weight multiplication product $\bm{H}$ enjoy sparsity from the previous input alternatively. In fact, as proven by (Zhang et al., 2023b), the scale of reduced computational operation can be advanced to be quadratic in $(1-\eta_{w})$.
5. Theoretical Analysis
In order to characterize Unifews for GNN learning, we firstly introduce a dedicated graph sparsifier for the approximate graph smoothing problem in Section 5.1. We subsequently examine the graph sparsification for decoupled GNN and the joint sparsification for iterative GNN in Sections 5.2 and 5.3, respectively. The insights derived from our following analysis are summarized as below:
  • Unifews can be described as a graph spectral sparsifier on both decoupled and iterative GNN architectures. Its sparsification strength is determined by the threshold of entry removal.
  • When applied across multiple GNN layers, Unifews provides an effective approximation to the graph smoothing optimization. Its output representation is close to the original objective.
  • Unifews offers inherent advantages in terms of GNN efficacy and efficiency, notably mitigating the over-smoothing issue and facilitating enhanced joint sparsity.
(a) Distribution of Entries
(b) Approximate Embedding Error
(c) Distance to Over-smoothed Embedding
Figure 3. Empirical evaluation on the SGC graph propagation and the effect of Unifews on cora dataset. (a) Distribution of entries in the diffusion matrix $\bm{T}$ and the message $\bm{\tau}_{(0)}$ at hop $l=0$. (b) The relative error margin $\|\hat{\bm{P}}_{(L)}-\bm{P}_{(L)}\|_{F}/\|\bm{P}_{(L)}\|_{F}$ of approximation to the raw embedding against the strength of Unifews and propagation hops. (c) The relative distance $\|\hat{\bm{P}}_{(L)}-\bm{P}^{\ast}\|_{F}/\|\bm{P}^{\ast}\|_{F}$ to the converged embedding against the strength of Unifews and propagation hops.
5.1. Approximation by Graph Sparsification
Graph sparsification methods such as edge removal apply modifications to the graph structure. We hereby look into a measure capable of quantitatively assessing the extent of graph changes induced by these sparsifiers. For this purpose, we introduce a variant of spectral similarity in light of its relevance to the overall GNN optimization goal. Consequently, our analysis highlights the impact of graph sparsifier throughout the learning process, rather than approximating specific feature values as in previous GNN sparsification theories (Srinivasa et al., 2020; Zhang et al., 2023b).
Definition 5.1 ($\epsilon$-Spectral Similarity).
The matrix $\hat{\bm{L}}$ is said to be $\epsilon$-spectrally similar to the matrix $\bm{L}$ if:
(13) $\bm{x}^{\top}(\hat{\bm{L}}-\epsilon\bm{I})\bm{x}\leq\bm{x}^{\top}\bm{L}\bm{x}\leq\bm{x}^{\top}(\hat{\bm{L}}+\epsilon\bm{I})\bm{x},\quad\forall\bm{x}\in\mathbb{R}^{n},$
or, equivalently:
(14) $\left|\bm{x}^{\top}(\bm{L}-\hat{\bm{L}})\bm{x}\right|\leq\epsilon\cdot\|\bm{x}\|^{2},\quad\forall\bm{x}\in\mathbb{R}^{n}.$
Spectral sparsification (Spielman and Srivastava, 2011; Batson et al., 2013) identifies a family of operations maintaining certain spectral properties such as eigenvalues of the graph during changes. Compared to the common multiplicative spectral similarity (Sadhanala et al., 2016; Calandriello et al., 2018; Charalambides and Hero, 2023), Definition 5.1 possesses an additive tolerance applied onto the graph Laplacian, which allows for modification on specific entries of the matrix $\bm{L}$ and suits our scenario.
Then we amend the graph Laplacian smoothing problem Definition 3.1 to conduct on the sparsified graph. If the corresponding Laplacian matrix adheres to the $\epsilon$-spectral similarity constraint in Definition 5.1, we have the following theorem for bounding the optimization goal of the approximate smoothing problem:
Theorem 5.2 (Approximate Graph Laplacian Smoothing).
Given two graphs $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ and $-^\hat{\mathcal{G}}=\langle\mathcal{V},\hat{\mathcal{E}}\rangle$, where $-^\hat{\mathcal{E}}$ is the sparsified edge set. The Laplacian matrix $\hat{\bm{L}}$ of $\hat{\mathcal{G}}$ is $\epsilon$-similar to $\bm{L}$ of $\mathcal{G}$. Then the solution $\hat{\bm{p}}^{\ast}$ to the problem Eq. 4 using $\hat{\bm{L}}$ is called an $\epsilon$-approximation of the solution $\bm{p}^{\ast}$ using $\bm{L}$ , and:
(15) $\|\hat{\bm{p}}^{\ast}-\bm{p}^{\ast}\|\leq c\epsilon\|\bm{p}^{\ast}\|.$
@@ 5.1. Approximation by Graph Sparsification
By using the closed-form solution in Eq. 5 and the fact that $\bm{A}^{-1}-\bm{B}^{-1}=\bm{B}^{-1}(\bm{B}-\bm{A})\bm{A}^{-1}$, we have:
@@
From Eq. 13, we can acquire the difference between the raw and approximate Laplacian matrices based on the spectral property:
(16) $\|\bm{L}-\hat{\bm{L}}\|_{2}=\sup_{\|\bm{x}\|=1}\bm{x}^{\top}(\bm{L}-\hat{\bm{L}})\bm{x}=\bm{x}_{0}^{\top}(\bm{L}-\hat{\bm{L}})\bm{x}_{0}\leq\epsilon\bm{x}_{0}^{\top}\bm{I}\bm{x}_{0}=\epsilon,$
@@
The distance between $\bm{p}^{\ast}$ and $\hat{\bm{p}}^{\ast}$ follows the consistency of spectral norm $\|\bm{Ax}\|\leq\|\bm{A}\|_{2}\|\bm{x}\|$. By substituting Eq. 16 and utilizing the property of spectral norm, we have:
@@
Given the additive nature of the graph specifier, the bound for graph smoothing problem Theorem 5.2 is dissimilar to the approximation setting with multiplicative similarity bounded by the quadratic form $O(c\epsilon\bm{p}^{\top}\bm{Lp})$ (Sadhanala et al., 2016; Calandriello et al., 2018), but instead correlates with the embedding vector norm $\|\bm{p}^{\ast}\|$. Theorem 5.2 establishes an interpretation for the smoothing procedure under a sparsified graph, that if a sparsifier complies with the spectral similarity Definition 5.1, it is capable of effectively approximating the iterative graph optimization and achieving a close output with bounded error.
5.2. Decoupled Case: Graph Sparsification
In this section, we primarily focus on Unifews applying to the decoupled propagation stage, where only graph sparsification Algorithm 1 takes effect. It provides a solitary scenario to inspect GNN propagation with the pruned graph.
We outline the diffusion in this case by the general graph adjacency $\bm{T}=\bm{A}$. The entry-wise difference matrix for the sparsified diffusion can be derived as $\bm{\Upsilon}=\bm{A}-\hat{\bm{A}}=\hat{\bm{L}}-\bm{L}$. Additionally, the pruned edges form the complement set $\mathcal{E}_{\Upsilon}=\mathcal{E}\setminus\hat{\mathcal{E}}$, and the number of removed edges is $q_{a}=|\mathcal{E}_{\Upsilon}|$. If an edge is pruned $(u,v)\in\mathcal{E}_{\Upsilon}$, the entry $\bm{\Upsilon}[u,v]=\bm{A}[u,v]$. By the following theorem, we show that the Unifews graph pruning for an arbitrary propagation layer can be considered as a spectral sparsification:
Theorem 5.3 ().
Given graph $\mathcal{G}$ and embedding $\bm{p}$, let $-^\hat{\mathcal{E}}$ be the edge set achieved by graph sparsification Eq. 8 with threshold $\delta_{a}$. The number of removed edges is $q_{a}$. Then, the sparsified Laplacian matrix $\hat{\bm{L}}$ is $\epsilon$-similar to $\bm{L}$ when $q_{a}\delta_{a}\leq\epsilon\|\bm{p}\|$.
@@ 5.2. Decoupled Case: Graph Sparsification
For a current embedding $\bm{p}$, its product with the difference matrix $\bm{\Upsilon p}$ only correlates with entries that have been pruned. Based on the Minkowski inequality and sparsification scheme Eq. 8, the $L_{1}$ norm of the product vector satisfies:
@@
Referring to Definition 5.1, when the condition $q_{a}\delta_{a}\leq\epsilon\|\bm{p}\|$ is met, the sparsified $\hat{\bm{L}}$ is spectrally similar to $\bm{L}$ with approximation rate denoted by $\epsilon$. ∎
The implication of Theorem 5.3 is that, the adaptive graph sparsification by Unifews for a single layer qualifies as a graph spectral sparsifier. The strength of sparsification represented by $\delta_{a}$, or equivalently, the relative threshold $\delta_{a}/\|\bm{p}\|$, determines the similarity to the raw graph matrix. As in Figure 2(a), a larger $\delta_{a}$ results in a greater quantity $q_{a}$ and consequently a further distance $\epsilon$.
@@
Practical GNNs are composed of multiple layers, which underscores the need of characterizing the sparsifier across layers. Consider consecutively sparsifying the diffusion for each hop by Eq. 10. Intuitively, as the edges are gradually removed from the original graph, there is $q_{a,(l_{1})}<q_{a,(l_{2})}$ for $l_{1}<l_{2}$ with the same relative threshold. If the sparsest graph $\bm{T}_{(L)}$ satisfies Definition 3.1, then the multi-layer update is also bounded. The following proposition states that, pruning and transmitting the sparsified graph still guarantees the approximation of the graph smoothing problem Theorem 5.2:
Proposition 5.4 ().
For an $L$-hop propagation with respective diffusion matrices $\bm{T}_{(l)}$, Algorithm 1 sparsification is applied to produce edge sets ^$\hat{\mathcal{E}}_{(l)}$. If the final sparsification in hop $L$ satisfies $\epsilon$-similarity, then the overall graph smoothing process is an $\epsilon$-approximation to the original problem.

### v2

@@ A.1 GNNs Simplification
Architecture Compression.
The concept of GNN pruning takes the neural network architecture into account and exploits sparsification for efficiency without hindering effectiveness (Zhang et al., 2023b). (Zhou et al., 2021) consider structural channel pruning by reducing the dimension of weight matrices, while CGP (Liu et al., 2023a) appends a prune-and-regrow scheme on the graph structure.
A line of research refers to the lottery ticket hypothesis (Frankle & Carbin, 2019; Hui et al., 2023) in the context of graph learning in search of a smaller yet effective substructure in the network. Among these approaches, GEBT (You et al., 2022) finds the initial subnetwork with simple magnitude-based pruning on both adjacency and weight matrices, whilst ICPG (Sui et al., 2024) learns the importance of edges as an external task. GLT (Chen et al., 2021) and DGLT (Wang et al., 2023b) employ an alternative optimization strategy that progressively processes the graph and network components. Other compression techniques such as quantization (Ding et al., 2021; Bahri et al., 2021; Wang et al., 2021b) are also investigated.
We mark the difference between our work and compression methods that the latter graph and network sparsification are loosely coupled, usually entailing a training procedure on the full-size graph and model to identify sparsification components, which prevents their application to large-scale data.
A.2 GNNs Modification
Graph Sampling.
Different from simplification, modification techniques change the message-passing scheme in GNNs, leading to new models with distinct expressiveness after modification. Sampling is a common strategy for improving GNN efficiency by altering the message-passing process. It is based on the intuition that the learning process only require a portion of data samples for each iteration. Thence, strategies on various hierarchies are designed for retrieving useful information (Chen et al., 2018; Huang et al., 2018; Zou et al., 2019; Zeng et al., 2020; Rong et al., 2020; Fang et al., 2023). A contemporary work (Ding et al., 2025) also investigates the utilization of spectral sparsification for sampling GNNs in the polynomial form, which echos with our theoretical analysis that graph sparsification process is useful in offering bounded approximation for message-passing GNNs. We however note that the sampling approach exhibits iterative processing of the graph components and does not produce a deterministic sparsified graph. Consequently, its capability on decoupled models, which are dominant on large graphs, is relatively constrained.
Graph Coarsening.
Graph coarsening describes another modification approach that reduces the graph size by contracting nodes into subsets. GraphZoom (Deng et al., 2020b), GOREN (Cai et al., 2021), and (Huang et al., 2021) explore various coarsening schemes and enhance training scalability, while GCond (Jin et al., 2022) alternatively adopts an analogous condensation strategy.
A.3 Comparative Discussion
@@
We highlight three advantages of our Unifews approach compared with above simplification and modification techniques in the context of GNN learning. Firstly, realistic graph propagation and feature transformation are conducted in an entry-centric fashion. Therefore, entry-level operations can be naturally inserted into the computation on the fly by modifying the matrix multiplication operator. Secondly, different from previous node- or edge-dependent strategies, our sparsification operates on the messages being passed during propagation, which is inherently informative to foster effective pruning while retaining model expressivity. Lastly, the progressive pruning across layers is capable of promoting greater graph sparsity, thereby alleviating intrinsic drawbacks of the GNN architecture including over-smoothing and neighborhood explosion. We empirically showcase the improvement on entry sparsity in Section C.3.
Appendix B Detailed Proof and Theoretical Analysis
B.1 Proof of Lemma 3.3
@@ Appendix B Detailed Proof and Theoretical Analysis
By using the closed-form solution $\bm{p}^{\ast}=(\bm{I}+c\bm{L})^{-1}\bm{x}$ and the fact that $\bm{A}^{-1}-\bm{B}^{-1}=\bm{B}^{-1}(\bm{B}-\bm{A})\bm{A}^{-1}$, we have:
@@
From Equation 2, we can acquire the difference between the raw and approximate Laplacian matrices based on the spectral property:
$\|\bm{L}-\hat{\bm{L}}\|_{2}=\sup_{\|\bm{x}\|=1}\bm{x}^{\top}(\bm{L}-\hat{\bm{L}})\bm{x}=\bm{x}_{0}^{\top}(\bm{L}-\hat{\bm{L}})\bm{x}_{0}\leq\epsilon\bm{x}_{0}^{\top}\bm{I}\bm{x}_{0}=\epsilon,$ (10)
@@
The distance between $\bm{p}^{\ast}$ and $\hat{\bm{p}}^{\ast}$ follows the consistency of spectral norm $\|\bm{Ax}\|\leq\|\bm{A}\|_{2}\|\bm{x}\|$. By substituting Equation 10 and utilizing the property of spectral norm, we have:
@@
Given the additive nature of the graph specifier, the bound for graph smoothing problem Lemma 3.3 is dissimilar to the approximation setting with multiplicative similarity bounded by the quadratic form $O(c\epsilon\bm{p}^{\top}\bm{Lp})$ (Sadhanala et al., 2016; Calandriello et al., 2018), but instead correlates with the embedding vector norm $\|\bm{p}^{\ast}\|$. This correlation arises from the bias introduced by the pruned entries in the diffusion matrix, which is associated with the embedding value.
B.2 Proof of Lemma 4.1
@@
We outline the diffusion by the general graph adjacency $\bm{T}=\bm{A}$. The entry-wise difference matrix for the sparsified diffusion can be derived as $\bm{\Upsilon}=\bm{A}-\hat{\bm{A}}=\hat{\bm{L}}-\bm{L}$. Additionally, the pruned edges form the complement set $\mathcal{E}_{\Upsilon}=\mathcal{E}\setminus\hat{\mathcal{E}}$, and the number of removed edges is $q_{a}=|\mathcal{E}_{\Upsilon}|$. If an edge is pruned $(u,v)\in\mathcal{E}_{\Upsilon}$, the entry $\bm{\Upsilon}[u,v]=\bm{A}[u,v]$.
For a current embedding $\bm{p}$, its product with the difference matrix $\bm{\Upsilon p}$ only correlates with entries that have been pruned. Based on the Minkowski inequality and sparsification scheme Equation 5, the $L_{1}$ norm of the product vector satisfies:
@@
Referring to Definition 3.2, Equation 2 is equivalent to:
$\left|\bm{x}^{\top}(\bm{L}-\hat{\bm{L}})\bm{x}\right|\leq\epsilon\cdot\|\bm{x}\|^{2},\quad\forall\bm{x}\in\mathbb{R}^{n}.$ (11)
@@
Therefore, when the condition $q_{a}\delta_{a}\leq\epsilon\|\bm{p}\|$ is met, the sparsified $\hat{\bm{L}}$ is spectrally similar to $\bm{L}$ with approximation rate $\epsilon$. ∎
B.3 Proof of Theorem 4.2
Proof.
Recall in Equation 5 pruning, there is $\mathcal{E}_{\Upsilon}=\{(u,v)\mid|\tau[u,v]|\leq\delta_{a}\}$. The fraction of entries below the threshold is thence expressed by the probability:
$q_{a}$ $=|\mathcal{E}_{\Upsilon}|=m\cdot P(\mathbf{A}[u,v]\cdot|\mathbf{p}[v]|\leq\delta_{a})$
[…]
where $f_{p[v]}(x)$ represents the probability distribution function of embedding values. As $\mathbf{A}[u,v]\geq x_{min}$, we have $P(\mathbf{A}[u,v]\leq\frac{\delta_{a}}{x})=0$, when $x>\frac{\delta_{a}}{x_{min}}$. The integral upper bound changes to $\frac{\delta_{a}}{x_{min}}$,i.e.
[…]
Acknowledging the assumptions, in scale-free graphs, the distribution of nodes follows the power law $P(d)=d^{-\alpha}$, where $P(d)$ is the fraction of nodes of degree $d$ for large values, and $\alpha$ is a constant typically within range $32<\alpha<$ 3. For the normalized adjacency matrix $\tilde{\bm{A}}=\bm{D}^{-1/2}\bar{\bm{A}}\bm{D}^{-1/2}$, its entry distribution also follows the power law as $x^{-\alpha}$. The embedding value $p[v]\sim N(0,\sigma^{2})$. By substituting the distributions we have:
[…]
$=2m\Bigl{[}\tfrac{1}{2}\,\mathrm{erf}\!\Bigl{(}\frac{\delta_{a}/x_{\min}}{\sqrt{2}\,\sigma}\Bigr{)}\;-\int_{0}^{\frac{\delta_{a}}{x_{min}}}\left(\frac{\delta_{a}}{xx_{min}}\right)^{-\alpha+1}\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{x^{2}}{2\sigma^{2}}\right)dx\bigl{]}$
[…]
$\text{where}\quad\mathrm{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}\mathrm{e}^{-t^{2}}\mathrm{d}t,\quad\gamma(s,x)=\int_{0}^{x}t^{\,s-1}\,\mathrm{e}^{-t}\,\mathrm{d}t.$
The last term can be given by the lower incomplete Gamma function:
[…]
Therefore,
[…]
To understand Equation 12, we discuss it in two extreme cases.
Case 1.
When $\delta_{a}\to 0$, only a small number of edges are affected. Considering the Taylor Expansion with respect to $\delta_{a}$, there is:
$q_{a}=c_{1}\,\delta+c_{3}\,\delta^{3}+\cdots,$
where $c_{1}=\frac{m\,\sqrt{2/\pi}}{\,x_{\min}\,\sigma}\,\frac{\alpha-1}{\alpha}$.
Case 2.
When $\delta_{a}>0$ is sufficiently large, this is the main case we concerned where a portion of edges are pruned. The formula will be approximated by:
$q_{a}=m-2m\frac{2^{\frac{\alpha-3}{2}}\Gamma\left(\frac{\alpha}{2}\right)}{\sqrt{\pi}\sigma^{\alpha+1}}\left(\frac{\delta_{a}}{x_{min}}\right)^{1-\alpha},$
or equivalently, $\eta_{a}=q_{a}/m=1-C\delta_{a}^{1-\alpha}$. Therefore, the relative strength of sparsification represented by the threshold $\delta_{a}/\|\bm{p}\|$ and the relative sparsity $\eta_{a}$ can be represented by each other.
Referring to Lemma 4.1, the approximation bound is given as:
$\epsilon=O\left(\eta_{a}(1-\eta_{a})^{\frac{1}{1-\alpha}}\right),$ (13)
which corresponds to Theorem 4.2 for giving the approximation bound. ∎
B.4 Proof Sketch of Proposition 4.3
@@
Consider consecutively sparsifying the diffusion for each hop by Equation 7. Intuitively, as the edges are gradually removed from the original graph, there is $q_{a,(l_{1})}<q_{a,(l_{2})}$ for $l_{1}<l_{2}$ with the same relative threshold. If the sparsest graph $\bm{T}_{(L)}$ satisfies Definition 3.1, then the multi-layer update is also bounded.